Porter Corporation makes and sells a single product called a Yute. The company is in the process of preparing its Selling and Administrative Expense Budget for the last quarter of the year. The following budget data are available:

Variable Cost
Per Yute Sold
Monthly Fixed
Cost
Sales commissions $ 5.90
Shipping $ 5.30
Advertising $ 8.90 $ 32,000
Executive salaries $ 178,000
Depreciation on office equipment $ 7,000
Other $ 0.60 $ 20,000

All of these expenses (except depreciation) are paid in cash in the month they are incurred.

If the company has budgeted to sell 14,000 Yutes in November, then the total budgeted selling and administrative expenses for November would be: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Variable cost per unit sold = (5.90+5.30+8.90+0.60) = 20.70

Fixed cost = 32000+178000+7000+20000 = 237000

Total budgeted selling and administrative = (14000*20.70+237000) = 526800